# A3VLM: Actionable Articulation-Aware Vision Language Model

### v1

@@ 4.1 Implementation Details
We fine-tuned the A3VLM model using the SPHINX framework [12] on eight NVIDIA A100 (80 GB) GPUs. The fine-tuning was completed in three epochs, which took approximately 24 hours. The visual encoders were kept frozen throughout the fine-tuning phase to maintain the integrity of the pre-trained features. We utilized the SPHINX-1K model, sourced directly from the official repository, as our pre-trained base. Training was conducted with a batch size of 4 and a learning rate set to $2\times 10^{-5}$. To effectively manage the computational load and enhance the training dynamics, we employed a gradient accumulation strategy with a factor of 4.
@@ 4.2 Qualitative Evaluation
To evaluate the action capabilities of A3VLM, we adopted the same settings as ManipLLM [21]. This benchmark utilizes Sapien [34] as the simulator and PartNet-Mobility [28] objects as the target objects. We use a flying Franka Panda Robot’s gripper with suction ability as the manipulator.
@@
Baselines. The robot manipulation success rates are compared for 20 training categories (used during training) and 10 testing categories (excluded from training). We compare our proposed method with five different baselines under the same setting: (1). Where2Act [27]: This method processes input from point-clouds to evaluate scores for each point. The highest-scoring point is chosen as the contact point. Additionally, it predicts 100 possible orientations for the end-effector, selecting the orientation with the top score to establish the contact pose. To ensure a fair comparison, we modified the originally used parallel gripper into a suction gripper. (2). UMPNet [36]: In alignment with UMPNet’s methodology, we carry out manipulations at the predicted contact point, positioning the end-effector’s orientation perpendicular to the surface of the object. (3). Flowbot3D [10]: This approach identifies the motion direction within a point cloud, referred to as “flow”. The point demonstrating the greatest flow magnitude is selected as the point of interaction, and the flow’s direction dictates the orientation of the end-effector. (4). Implicit3D [40]: This framework formulates a manipulation strategy for subsequent tasks by employing the Transporter network to identify keypoints on 3D articulated objects. These keypoints facilitate the determination of the end-effector’s pose. (5). ManipLLM [21]: ManipLLM is the current state-of-the-art method on this benchmark. It uses a vision language model to predict the contact point and forward direction of a suction gripper given an RGBD image and a language prompt.
Result. Table. 2 shows the performance of A3VLM and five other baselines. From the table, we can see that A3VLM outperforms all baselines by a large margin for most object categories. This result demonstrates the effectiveness of A3VLM for manipulating articulated objects.

### v2

@@ 4.1 Implementation Details
We fine-tuned the A3VLM model using the SPHINX framework [12] on eight NVIDIA A100 (80 GB) GPUs. The fine-tuning was completed in three epochs, which took approximately 24 hours. The visual encoders were kept frozen throughout the fine-tuning phase to maintain the integrity of the pre-trained features. We utilized the SPHINX-1K model, sourced directly from the official repository, as our pre-trained base. Training was conducted with a batch size of 4 and a learning rate set to $2\times 10^{-5}$.
@@ 4.2 Qualitative Evaluation
To evaluate the action capabilities of A3VLM, we modified upon the settings of ManipLLM [21]. This benchmark utilizes Sapien [34] as the simulator and PartNet-Mobility [28] objects as the target objects. Object’s joint values will be initialized to the middle value. We use a flying Franka Panda Robot’s gripper with suction ability as the manipulator.
@@
Baselines. The robot manipulation success rates are compared for 20 training categories (used during training) and 10 testing categories (excluded from training). We compare our proposed method with five different baselines: (1). Where2Act [27]: This method processes input from point-clouds to evaluate scores for each point. The highest-scoring point is chosen as the contact point. Additionally, it predicts 100 possible orientations for the end-effector, selecting the orientation with the top score to establish the contact pose. To ensure a fair comparison, we modified the originally used parallel gripper into a suction gripper. (2). UMPNet [36]: In alignment with UMPNet’s methodology, we carry out manipulations at the predicted contact point, positioning the end-effector’s orientation perpendicular to the surface of the object. (3). Flowbot3D [10]: This approach identifies the motion direction within a point cloud, referred to as “flow”. The point demonstrating the greatest flow magnitude is selected as the point of interaction, and the flow’s direction dictates the orientation of the end-effector. (4). Implicit3D [40]: This framework formulates a manipulation strategy for subsequent tasks by employing the Transporter network to identify keypoints on 3D articulated objects. These keypoints facilitate the determination of the end-effector’s pose. (5). ManipLLM [21]: ManipLLM is the current state-of-the-art method on this benchmark. It uses a vision language model to predict the contact point and forward direction of a suction gripper given an RGBD image and a language prompt.
Action Primitives Details. Different from the mentioned baselines, A3VLM models the action in an object-centric way. To be more specific, for each object, we first detect a list of action parts with corresponding bounding boxes $\mathcal{B}$, axes $\mathcal{A}$, joint types, and link names $\mathcal{S}$. We select a random action part from the list and use its bounding box $\mathcal{B}$ and axis $\mathcal{A}$ to generate two robot trajectories. For example, for the handle of a faucet, we will generate trajectories to rotate it clockwise and counter-clockwise. We will execute these trajectories in two attempts. The task is regarded as successful if it succeeds in either try.
Result. Table. 2 shows the performance of A3VLM and five other baselines. From the table, we can see that A3VLM outperforms all baselines by a large margin for most object categories. We think this improvement comes from two folds. One fold comes from the accurate grounding of actionable parts and articulation structure and the other fold is the introduce of action primitives. Action primitives enable A3VLM to perform different actions towards different articulated objects. This result demonstrates the effectiveness of A3VLM for manipulating articulated objects.